OpenStudy (anonymous):

all pairs (a,b) of positive integer satisfying : (a^2 + b)|(a^2*b+a) and (b^2 - a)|(ab^2 + b)

OpenStudy (anonymous):

i guess to solve one by one for (a^2 + b)|(a^2*b+a) and (b^2 - a)|(ab^2 + b) then find intersect, right ?

OpenStudy (loser66):

is it not that the condition is \(a^2b+a\) is divided by \(a^2+b\)??

OpenStudy (loser66):

dan interpreted the problem in another way, right?

OpenStudy (dan815):

i dunno man im just moving stuff aroudn

OpenStudy (dan815):

u cud do that too i guess, and do some sythetic division

OpenStudy (anonymous):

yeah, in other words a^2 + b is divisible by a^2 * b + a and .... @Loser66

OpenStudy (dan815):

ya u are right

OpenStudy (dan815):

a * b = c * d a b c d are all integers so

OpenStudy (dan815):

but it doesnt mean they have to be divisible

OpenStudy (anonymous):

ah , if i use the long division get a = b^2 for the first one

OpenStudy (dan815):

like (p1*p2) * (p3*p4) = (p3*p2*p1) * (p4) this could happen and p1*p2 is not divisible by p4

OpenStudy (rational):

is b = a+1 the ansur ?

OpenStudy (anonymous):

how ? btw it is not options question, so i dont know

OpenStudy (rational):

it seem to work, check once

OpenStudy (rational):

I'll put the solution shortly..

OpenStudy (dan815):

omg i didnt see that Division sign thing

OpenStudy (anonymous):

give me a little hint

OpenStudy (rational):

lmao

OpenStudy (dan815):

thats hard to see on my computer lol

OpenStudy (rational):

can you double check if it is really working or not i don't have a rigorous proof yet, still working..

OpenStudy (dan815):

so we are basically solving a system of modular equatiosn right

OpenStudy (anonymous):

yes, b = a + 1 is satisfying both for (a^2 + b)|(a^2*b+a) and (b^2 - a)|(ab^2 + b)

OpenStudy (dan815):